OpenStudy (anonymous):

A cake recipe calls for one and two third cups of flour. If Gloria is planning on making 6 cakes for a bake sale, how many cups of flour will she need?

OpenStudy (gorv):

1cake =1+2/3 cup 6 cake = 6*1+6*2/3

OpenStudy (gorv):

=6+2*2=6+4=10

OpenStudy (gorv):

welcome to OS :)

OpenStudy (gorv):

aske OS related queries here http://openstudy.com/study#/groups/OpenStudy%20Ambassadors